# Models of interacting bosons with exact ground states: a unified approach

### v2

@@ Abstract
Dipole conserving dynamics: It is sometimes interesting to consider Hamiltonians with dynamics that preserve the center of mass position. Models of this sort can arise in the low energy effective theory of strongly interacting particles where remaining dynamical processes must satisfy strong constraints. Examples of this are the ring-exchange processes in quantum crystals Thouless (1965), the “resonance” terms in the QDM Kivelson et al. (1987), induced pair-hopping terms in quantum Hall systems in the thin cylinder limit Bergholtz and Karlhede (2006, 2008); Seidel et al. (2005), and the hopping processes in a strongly tilted optical lattice Khemani et al. (2020).
@@
Many aspects of systems with this sort of kinetic energy can be treated from the present perspective by suitably defining the clusters $c$ and $c^{\prime}$ connected by $\hat{H}_{c,c^{\prime}}$. The resulting solvable ground states can still be compressible and possess ODLRO. However, for processes that conserve the center-of-mass position of the bosons, $g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})$ vanishes at least as $\bm{q}^{2}$ as $\bm{q}\to\bm{0}$, implying that $f(\bm{q})\sim|\bm{q}|^{4}$ as $\bm{q}\to\bm{0}$ (see Eq. 21). It follows from Eq. 24 that the superfluid stiffness (i.e. the coefficient of $\left({\bf\nabla}\theta\right)^{2}$ in the effective action) vanishes. Since any uniform current-carrying state would violate the assumed dipole conservation, such states are also, in some sense, insulating - they have thus been named ‘Bose-Einstein insulators Lake et al. (2022)’.
@@
Bose liquids with ‘Bose surfaces’: When the non-zero $\tau_{c,c^{\prime}}$’s all correspond to $\left|g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})\right|^{2}$ that vanish on a submanifold $\mathcal{M}$ of the momentum space, this implies that $f(\bm{q})$ as well as $E_{\text{SMA}}(\bm{q})$ (assuming the system is compressible) vanish on $\mathcal{M}$. For most cases, $\mathcal{M}$ is a single point at $\bm{q}=\bm{0}$, but sometimes it can be higher dimensional. As a concrete example, consider short-range resonant pair hopping on a square lattice where a pair of bosons from one set of opposite vertices, $c$, of an elementary plaquette is shifted to the complementary vertices, $c^{\prime}$. [Note that this process happens to be dipole conserving; moreover, if we identify the bosons with dimers and the lattice vertices with the edge centers of another rotated square lattice, this is the sort of dimer resonance term mentioned above that must be introduced to obtain the dynamical terms in the QDM at $\bar{N}=1$.] The resulting $f(\bm{q})\sim\sin^{2}q_{x}\sin^{2}q_{y}$ not only vanishes as $\bm{q}\rightarrow\bm{0}$, but also along lines in $\bm{q}$-space. Therefore, the system has gapless modes along one-dimensional ‘Bose-surfaces’ Motrunich and Fisher (2007) in the Brillouin zone.
@@
The addition of a density interaction $w(\delta n)^{2}/2$ to the effective action makes the system unstable to phase separation (as usual) for attractive $w<0$. In contrast, for repulsive $w>0$, it keeps the gapless manifold $\mathcal{M}$ but changes the boson excitation dispersion near $\mathcal{M}$ from quadratic to linear: $E(\bm{q})\sim\sqrt{wf(\bm{q})}$, which results in a spatially decaying boson Green’s function $\propto\langle\mathrm{e}^{\mathrm{i}[\theta(\bm{r})-\theta(\bm{0})]}\rangle$ with a similar character as for a Fermi liquid with the same manifold $\mathcal{M}$ of gapless modes (e.g. a surface in 3D or a nodal line in 2D)  [^6] It should be noted that the state constructed here is different from the Bose liquid studied in Ref. Motrunich and Fisher (2007) in many aspects, including its specific heat and the relation between the particle density and the volume enclosed by the Bose surfaces. .
@@
Discussion: In this work, we have developed a unified approach for constructing “frustration-free” Hamiltonians for number-conserving bosons whose exact ground-state properties can be inferred from the properties of an associated classical statistical mechanical problem on the same lattice. For the known solvable models, our scheme provides not only an alternative representation that might be realizable on engineered platforms like optical lattices but also a natural route from a strongly constrained problem to a free (trivial) limit of the boson problem. When the constrained limit of the problem exhibits an exotic quantum phase of matter, this evolution must drive exotic transitions out of them, whose properties may be studied with these engineered models. The critical points constructed this way are “conformal quantum critical points” Ardonne et al. (2004) with unusual dynamical properties Castelnovo et al. (2005); Isakov et al. (2011).
@@
The asymptotic behaviors of correlations in $d$ dimensional quantum ($T=0$) and classical ($T>0$) systems are generically different. Thus, solvable models of the sort we have discussed are always, to some degree, non-generic. Indeed, in many cases, as we have shown, the solvable models describe systems on the phase boundary beyond which there is instability - often to phase separation. Still, it is possible to infer generic properties of interesting phases by constructing an effective field theoretical representation of the solvable model and then addressing the effect of symmetry-allowed perturbations.
@@
There are a variety of further directions that presumably can be successfully explored by extensions of the present approach. For one thing, disordered versions of the present models can easily be constructed: An artificial version of this (similar to the Mattis model Mattis (1976)) is achieved by choosing random hopping matrix elements, $\tau_{c,c^{\prime}}$, while keeping $K$ translationally invariant, which results in a translationally invariant ground state, but where the disorder is strongly reflected in the excitation spectrum. More interestingly, if $K$ is not translationally invariant, then the resulting potential energy will not be either, nor will the associated ground state, despite still exhibiting ODLRO.
@@
Acknowledgements: We thank Jonah Herzog-Arbeitman, Steven Girvin, Vladimir Calvera, Kyung-Su Kim, and Eduardo Fradkin for helpful communications. This work was supported, in part, by the U. S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Award No. DE-SC0020045.

### v3

@@ Abstract
Dipole conserving dynamics: It is sometimes interesting to consider Hamiltonians with dynamics that preserve the center of mass position. Models of this sort can arise in the low energy effective theory of strongly interacting particles where remaining dynamical processes must satisfy strong constraints. Examples of this are the ring-exchange processes in quantum crystals [88], the “resonance” terms in the QDM [6], induced pair-hopping terms in quantum Hall systems in the thin cylinder limit [89, 90, 91], and the hopping processes in a strongly tilted optical lattice [92].
@@
Many aspects of systems with this sort of kinetic energy can be treated from the present perspective by suitably defining the clusters $c$ and $c^{\prime}$ connected by $\hat{H}_{c,c^{\prime}}$. The resulting solvable ground states can still be compressible and possess ODLRO. However, for processes that conserve the center-of-mass position of the bosons, $g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})$ vanishes at least as $\bm{q}^{2}$ as $\bm{q}\to\bm{0}$, implying that $f(\bm{q})\sim|\bm{q}|^{4}$ as $\bm{q}\to\bm{0}$ (see Eq. 18). It follows from Eq. 20 that the superfluid stiffness (i.e. the coefficient of $\left({\bf\nabla}\theta\right)^{2}$ in the effective action) vanishes. Since any uniform current-carrying state would violate the assumed dipole conservation, such states are also, in some sense, insulating - they have thus been named ‘Bose-Einstein insulators [93]’.
@@
Bose liquids with ‘Bose surfaces’: When the non-zero $\tau_{c,c^{\prime}}$’s all correspond to $\left|g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})\right|^{2}$ that vanish on a submanifold $\mathcal{M}$ of the momentum space, this implies that $f(\bm{q})$ as well as $E_{\text{SMA}}(\bm{q})$ (assuming the system is compressible) vanish on $\mathcal{M}$. For most cases, $\mathcal{M}$ is a single point at $\bm{q}=\bm{0}$, but sometimes it can be higher dimensional. As a concrete example, consider short-range resonant pair hopping on a square lattice where a pair of bosons from one set of opposite vertices, $c$, of an elementary plaquette is shifted to the complementary vertices, $c^{\prime}$. [Note that this process happens to be dipole conserving; moreover, if we identify the bosons with dimers and the lattice vertices with the edge centers of another rotated square lattice, this is the sort of dimer resonance term mentioned above that must be introduced to obtain the dynamical terms in the QDM at $\bar{N}=1$.] The resulting $f(\bm{q})\sim\sin^{2}q_{x}\sin^{2}q_{y}$ not only vanishes as $\bm{q}\rightarrow\bm{0}$, but also along lines in $\bm{q}$-space. Therefore, the system has gapless modes along one-dimensional ‘Bose-surfaces’ [94] in the Brillouin zone.
@@
The addition of a density interaction $w(\delta n)^{2}/2$ to the effective action makes the system unstable to phase separation (as usual) for attractive $w<0$. In contrast, for repulsive $w>0$, it keeps the gapless manifold $\mathcal{M}$ but changes the boson excitation dispersion near $\mathcal{M}$ from quadratic to linear: $E(\bm{q})\sim\sqrt{wf(\bm{q})}$, which results in a spatially decaying boson Green’s function $\propto\langle\mathrm{e}^{\mathrm{i}[\theta(\bm{r})-\theta(\bm{0})]}\rangle$ with a similar character as for a Fermi liquid with the same manifold $\mathcal{M}$ of gapless modes (e.g. a surface in 3D or a nodal line in 2D)  [^5] It should be noted that the state constructed here is different from the Bose liquid studied in Ref. [94] in many aspects, including its specific heat and the relation between the particle density and the volume enclosed by the Bose surfaces. .
@@
Discussion: In this work, we have developed a unified approach for constructing “frustration-free” Hamiltonians for number-conserving bosons whose exact ground-state properties can be inferred from the properties of an associated classical statistical mechanical problem on the same lattice. For the known solvable models, our scheme provides not only an alternative representation that might be realizable on engineered platforms like optical lattices but also a natural route from a strongly constrained problem to a free (trivial) limit of the boson problem. When the constrained limit of the problem exhibits an exotic quantum phase of matter, this evolution must drive exotic transitions out of them, whose properties may be studied with these engineered models. The critical points constructed this way are “conformal quantum critical points” [13] with unusual dynamical properties [16, 96].
@@
The asymptotic behaviors of correlations in $d$ dimensional quantum ($T=0$) and classical ($T>0$) systems are generically different. Thus, solvable models of the sort we have discussed are always, to some degree, non-generic. Indeed, in many cases, as we have shown, the solvable models describe systems on the phase boundary beyond which there is instability - often to phase separation. A dramatic manifestation of this point is that the charge $U(1)$ symmetry is broken in the ground state of such models even in one dimension, similar to certain frustration-free spin models [97]. Still, it is possible to infer generic properties of interesting phases by constructing an effective field theoretical representation of the solvable model and then addressing the effect of symmetry-allowed perturbations.
@@
There are a variety of further directions that presumably can be successfully explored by extensions of the present approach. For one thing, disordered versions of the present models can easily be constructed: An artificial version of this (similar to the Mattis model [98]) is achieved by choosing random hopping matrix elements, $\tau_{c,c^{\prime}}$, while keeping $K$ translationally invariant, which results in a translationally invariant ground state, but where the disorder is strongly reflected in the excitation spectrum. More interestingly, if $K$ is not translationally invariant, then the resulting potential energy will not be either, nor will the associated ground state, despite still exhibiting ODLRO.
@@
Acknowledgements: We thank Jonah Herzog-Arbeitman, Steven Girvin, Vladimir Calvera, Kyung-Su Kim, Haruki Watanabe, Biao Lian, and Eduardo Fradkin for helpful communications. This work was supported, in part, by the U. S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, under Award No. DE-SC0020045.
Appendix A Generalized Witten’s conjugation method
Here, we formally discuss the construction scheme we adopted in this work, which could be viewed as a generalization of Witten’s conjugation method [34, 70].
Suppose we are given a frustration-free Hamiltonian
$\hat{H}=\sum_{l}\hat{\mathcal{O}}_{l}^{\dagger}\hat{\mathcal{O}}_{l}$ (26)
and its ground state, $|\Psi_{0}\rangle$, such that $\hat{\mathcal{O}}_{l}|\Psi_{0}\rangle=0$. Then one can construct new frustration-free models of the form
$\hat{H}=\sum_{l}\hat{\mathcal{P}}^{-1,\dagger}\hat{\mathcal{O}}_{l}^{\dagger}\hat{\mathcal{M}}_{l}\hat{\mathcal{O}}_{l}\hat{\mathcal{P}}^{-1}$ (27)
with ground state $|\Psi\rangle\equiv\hat{\mathcal{P}}|\Psi_{0}\rangle$, where $\hat{\mathcal{P}}$ is an arbitrary invertible operator and {^$\{\hat{\mathcal{M}}_{l}\}$ is a set of arbitrary positive semi-definite operators. This is because each term in $\hat{H}$ is still positive semi-definite and annihilates $|\Psi\rangle$.
The specific class of models discussed in the main text is of this form with $\hat{\mathcal{O}}_{l}\rightarrow\hat{\mathcal{B}}_{c,c^{\prime}}$, $\hat{\mathcal{P}}\rightarrow\hat{P}$, and ^$\hat{\mathcal{M}}_{l}\rightarrow\hat{P}_{c}\hat{P}_{c^{\prime}}$, the later operator so chosen as to ensure that the resulting hopping terms retains a simple and physical form. It is also similar to a construction for interacting boson models in the continuum studied in Ref. [14] where the engineered Hamiltonian is written as
$\hat{\mathcal{H}}=$ $\hat{P}^{-1}\left(\nabla\psi^{\dagger}\right)\hat{P}\cdot\hat{P}\left(\nabla\psi\right)\hat{P}^{-1}.$ (28)
with a similar $\hat{P}$ related to the Boltzmann weight of a classical gas problem in the continuum.
Appendix B The number basis
Since all the $|\alpha\rangle$ have zero energy, any linear combination of them is also a ground state of the system. Since the models are number-conserving, we can obtain a number basis for the ground state manifold:
$\left|M\right\rangle\equiv$ $\frac{1}{\mathcal{N}_{M}}\ \hat{P}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}\left|0\right\rangle$ (29)
$=$ $\frac{M!\mathcal{N}_{|\alpha_{0}|}}{|\alpha|^{M}\mathcal{N}_{M}}\int_{0}^{2\pi}\frac{\mathrm{d}\theta}{2\pi}\mathrm{e}^{-\mathrm{i}M\theta}\left|\alpha=|\alpha_{0}|\mathrm{e}^{\mathrm{i}\theta}\right\rangle\ .$ (30)
which is true for any fixed non-zero value of $|\alpha_{0}|$.
As long as $\tau_{c,c^{\prime}}$ is generic enough for the kinetic Hamiltonian to connect all the possible configuration states |$|\{n\}\rangle$ within the corresponding number sector, the ground state $|M\rangle$ is unique (up to potential degeneracies associated with spontaneously broken symmetries or topological order) due to the Perron-Frobenius theorem, since the off-diagonal (kinetic) part of the Hamiltonian in this basis is strictly non-positive and irreducible.
$\mathcal{N}_{M}$ can be evaluated as:
$\mathcal{N}_{M}^{2}=$ $\left\langle 0\right|\left(\sum_{i}\hat{b}_{i}\right)^{M}\hat{P}^{2}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}\left|0\right\rangle$ (31)
$=$ $\sum_{\{n\}|\sum_{k}n_{k}=M}{M\choose n_{1},\dots,n_{\mathsf{V}}}^{2}\mathrm{e}^{-K(\{n\})}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \langle 0|\prod_{k}(\hat{b}_{k})^{n_{k}}(\hat{b}_{k}^{\dagger})^{n_{k}}|0\rangle$ (32)
$=$ $(M!)^{2}\cdot\sum_{\{n\}|\sum_{k}n_{k}=M}\frac{1}{\prod_{k}(n_{k}!)}\mathrm{e}^{-K(\{n\})}$ (33)
$\equiv$ $(M!)^{2}\mathcal{Z}_{M}$ (34)
where $\mathcal{Z}_{M}$ coincides with the partition function of a classical lattice gas problem in its canonical ensemble. The problem is defined by putting $M$ indistinguishable particles on the lattice sites, assigning each occupation configuration $\{n\}$ an interaction ‘energy’ $K(\{n\})$, and lastly, setting the effective ‘temperature’ of the classical gas to $1$.
Appendix C The mapping to classical problem
Since
$\left|\alpha\right\rangle=$ $\frac{1}{\mathcal{N}_{\alpha}}\hat{P}\sum_{M=0}^{\infty}\exp\left[\alpha\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)\right]\left|0\right\rangle$ (35)
$=$ $\frac{1}{\mathcal{N}_{\alpha}}\hat{P}\sum_{M=0}^{\infty}\frac{\alpha^{M}}{M!}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}\left|0\right\rangle$ (36)
$=$ $\frac{1}{\mathcal{N}_{\alpha}}\sum_{M=0}^{\infty}\frac{\alpha^{M}\mathcal{N}_{M}|M\rangle}{M!}$ (37)
$\mathcal{N}_{\alpha}$ can be evaluated as:
$\mathcal{N}^{2}_{\alpha}=$ $\sum_{M}\frac{|\alpha|^{2M}}{(M!)^{2}}\mathcal{N}^{2}_{M}$ (38)
$=$ $\sum_{\{n\}}\frac{1}{\prod_{k}(n_{k}!)}\mathrm{e}^{-K(\{n\})+\mu_{\text{cl}}\sum_{k}n_{k}}$ (39)
$\equiv$ $\mathcal{Z}_{\alpha}$ (40)
where the summation over occupation configurations $\{n\}$ is now unrestricted, and $\mathcal{Z}_{\alpha}$ can be understood as the partition function of the above classical lattice gas problem in the grand canonical ensemble with $\mu_{\text{cl}}\equiv\ln(|\alpha|^{2})$ being the chemical potential. The total particle number expectation can be thus evaluated as
$\langle\alpha|\sum_{i}\hat{n}_{i}|\alpha\rangle=\frac{\partial\ln\mathcal{Z}_{\alpha}}{\partial\mu_{\text{cl}}}$ (41)
It is straightforward to show that all density correlations in the quantum ground state agree with those in the classical system. We now further show that the boson correlation functions can also be conveniently calculated as with density correlators in the classical system (here $c$ and $c^{\prime}$ can be of different sizes, $L$ and $L^{\prime}$):
$\left\langle M\right|\hat{B}^{\dagger}_{c}\hat{B}_{c^{\prime}}\left|M\right\rangle=$ $\delta_{L,L^{\prime}}\sum_{\{n\}|\sum_{k}n_{k}=M-L}\left\langle M\right|\hat{B}^{\dagger}_{c}|\{n\}\rangle\langle\{n\}|\hat{B}_{c^{\prime}}\left|M\right\rangle$ (42)
$=$ $\frac{\delta_{L,L^{\prime}}}{\mathcal{N}_{M}}\sum_{\{n\}|\sum_{k}n_{k}=M-L}\frac{1}{\prod_{k}(n_{k}!)}\langle 0|\left(\sum_{i}\hat{b}_{i}\right)^{M}\hat{P}\hat{B}^{\dagger}_{c}\prod_{k}(\hat{b}^{\dagger}_{k})^{n_{k}}|0\rangle\langle 0|\prod_{k}(\hat{b}_{k})^{n_{k}}\hat{B}_{c^{\prime}}\hat{P}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}|0\rangle$ (43)
$=$ $\frac{\delta_{L,L^{\prime}}}{\mathcal{N}_{M}}\sum_{\{n\}|\sum_{k}n_{k}=M-L}\frac{(M!)^{2}}{\prod_{k}(n_{k}!)}\mathrm{e}^{-\delta K_{c}(\{n\})/2-\delta K_{c^{\prime}}(\{n\})/2}\mathrm{e}^{-K(\{n\})}$ (44)
$=$ $\frac{\delta_{L,L^{\prime}}\mathcal{Z}_{M-L}}{\mathcal{Z}_{M}}\Big{\langle}\!\!\Big{\langle}\mathrm{e}^{-\delta K_{i_{1}\dots i_{L}}(\{n\})/2}\mathrm{e}^{-\delta K_{j_{1}\dots j_{L}}(\{n\})/2}\Big{\rangle}\!\!\Big{\rangle}_{M-L}$ (45)
$\left\langle\alpha\right|\hat{B}^{\dagger}_{i_{1}\dots i_{L_{+}}}\hat{B}_{j_{1}\dots j_{L_{-}}}\left|\alpha\right\rangle=$$\sum_{\{n\}}\langle\alpha|\hat{B}^{\dagger}_{i_{1}\dots i_{L_{+}}}|\{n\}\rangle\langle\{n\}|\hat{B}_{j_{1}\dots j_{L_{-}}}|\alpha\rangle$ (46)
$=$ $\frac{(\alpha^{*})^{L_{+}}\alpha^{L_{-}}}{\mathcal{Z}_{\alpha}}\sum_{\{n\}}\frac{1}{\prod_{k}(n_{k}!)}\mathrm{e}^{-\delta K_{c}(\{n\})/2-\delta K_{c^{\prime}}(\{n\})/2}\mathrm{e}^{-K(\{n\})}$ (47)
$=$ $(\alpha^{*})^{L_{+}}\alpha^{L_{-}}\Big{\langle}\!\!\Big{\langle}\mathrm{e}^{-\delta K_{c}(\{n\})/2}\mathrm{e}^{-\delta K_{c^{\prime}}(\{n\})/2}\Big{\rangle}\!\!\Big{\rangle}_{\alpha}$ (48)
Appendix D Single mode approximation
The variational energy
$E_{\text{SMA}}(\bm{q})=\frac{\left\langle\alpha\left|\hat{\rho}_{\bm{q}}^{\dagger}\hat{H}\hat{\rho}_{\bm{q}}\right|\alpha\right\rangle}{\left\langle\alpha\left|\hat{\rho}_{\bm{q}}^{\dagger}\hat{\rho}_{\bm{q}}\right|\alpha\right\rangle}$ (49)
By definition, the denominator is $s(\bm{q})\mathsf{V}$. Then, using $\hat{H}|\alpha\rangle=0$, we can rewrite the numerator as
$\mathsf{V}f(\bm{q})=\left\langle\alpha\left|\hat{\rho}_{\bm{q}}^{\dagger}[\hat{H},\hat{\rho}_{\bm{q}}]\right|\alpha\right\rangle=\left\langle\alpha\left|\hat{\rho}_{\bm{q}}^{\dagger}\partial_{\tau}\hat{\rho}_{\bm{q}}(\tau)|_{\tau=0}\right|\alpha\right\rangle$ (50)
which is proportional to the time derivative of the dynamical structure factor at equal time and hence its first frequency moment. Due to inversion symmetry we assumed, $f(\bm{q})=f(-\bm{q})=f^{*}(\bm{q})$, so
$f(\bm{q})=$ $\frac{f(\bm{q})+f(-\bm{q})}{2}$ (51)
$=$ $\frac{1}{2\mathsf{V}}\left\langle\alpha\left|[\hat{\rho}_{\bm{q}}^{\dagger},[\hat{H},\hat{\rho}_{\bm{q}}]]\right|\alpha\right\rangle$ (52)
Evaluating the double commutator gives rise to the expression in Eq. 18.
We note that the tightness of the upper bound obtained this way can be further estimated by the variance
$\left\langle\left(\hat{H}-E_{\text{SMA}}(\bm{q})\right)^{2}\right\rangle=\frac{\left\langle M\right|[\hat{\rho}_{\bm{q}}^{\dagger},\hat{H}][\hat{H},\hat{\rho}_{\bm{q}}]\left|M\right\rangle}{\left\langle M\right|\hat{\rho}_{\bm{q}}^{\dagger}\hat{\rho}_{\bm{q}}\left|M\right\rangle}-E^{2}_{\text{SMA}}(\bm{q})$ (53)
Appendix E Generalization to number non-conserving systems
One can generalize the current construction scheme to number non-conserving systems in several ways. The first is to allow the $c,c^{\prime}$ in Eq. 8 to have different sizes, and modify $\left(\hat{B}_{c}-\hat{B}_{c^{\prime}}\right)$ to
$\left(\hat{B}_{c}/\alpha_{0}^{L_{c}}-\hat{B}_{c^{\prime}}/\alpha_{0}^{L_{c^{\prime}}}\right)$ (54)
where $L_{c}$ is the size of the cluster $c$. The $\hat{H}_{c,c^{\prime}}$ constructed in this way does not conserve particle number, and only certain modified coherent states $|\alpha=\alpha_{0}\mathrm{e}^{\mathrm{i}2\pi r/q}\rangle$ can be shown to be the ground state, where $q=\text{gcd}(L_{c},L_{c^{\prime}})$ and $r=0,1,\dots q-1$.
A further generalization could be made by defining modified coherent states with a set of site-dependent $\{\alpha_{i}\}$:
|$\left|\{\alpha_{i}\}\right\rangle\equiv$ $\frac{1}{\mathcal{N}_{\{\alpha_{i}\}}}\ \hat{P}\exp\left[\sum_{i}\alpha_{i}\hat{b}^{\dagger}_{i}\right]\left|0\right\rangle\ .$ (55)
then further modifying $\left(\hat{B}_{c}-\hat{B}_{c^{\prime}}\right)$ in Eq. 8 to
$\left[\hat{B}_{c}/(\prod_{j\in c}\alpha_{0,j})-\hat{B}_{c^{\prime}}/(\prod_{j\in c^{\prime}}\alpha_{0,j})\right]$ (56)
will give rise to the parent Hamiltonians for ground states |$\left|\{\alpha_{i}=\alpha_{0,i}\mathrm{e}^{\mathrm{i}2\pi r/q}\}\right\rangle$.